OpenStudy (anonymous):

f(g(1/2) f(x)=4x+5 g(x)=sqrt 5

OpenStudy (cggurumanjunath):

u try i will help you !

OpenStudy (anonymous):

sqrt x sorry

OpenStudy (cggurumanjunath):

first find g(1/2).

OpenStudy (cggurumanjunath):

@ninab731

OpenStudy (anonymous):

\[f(g(\frac{ 1 }{ 2 }))\] \[\sqrt{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

\[g(1/2) = \sqrt{1/2}\] So now you can put that into \[f(x) = 4x+5\]

OpenStudy (cggurumanjunath):

now find f(1/sqrt 2)

OpenStudy (cggurumanjunath):

@ninab731

OpenStudy (cggurumanjunath):

substitute x=1/sqrt 2 in 4x+5

OpenStudy (anonymous):

get some weird decimal answer says \[5+2\sqrt{2}\]

OpenStudy (cggurumanjunath):

u got it :)

OpenStudy (cggurumanjunath):

@ninab731

OpenStudy (cggurumanjunath):

well done .

OpenStudy (anonymous):

i get some decimal number not that answer is from book

OpenStudy (cggurumanjunath):

use a calculator then !

OpenStudy (cggurumanjunath):

7.828

OpenStudy (anonymous):

ya thats what i get but how do they get it into the form of \[5+2\sqrt{2}\]

OpenStudy (vjordillo):

You have the same answer to your book. You just got the decimal equivalence of 2 sqrt of 2

zepdrix (zepdrix):

\[\Large g(x)=\sqrt x\]\[\Large \color{royalblue}{g(1/2)=\sqrt {1/2}}\] \[\Large f(\color{royalblue}{x})\quad=\quad 4(\color{royalblue}{x})+5\]\[\Large f(\color{royalblue}{g(1/2)})\quad=\quad 4(\color{royalblue}{\sqrt {1/2}})+5\]

OpenStudy (cggurumanjunath):

tell me sqrt 2 use a calc.

OpenStudy (cggurumanjunath):

do u have a calculator at first @ninab731 ?

OpenStudy (vjordillo):

And 2sqrt of 2is 2.828. Adding its value to 5, makes your answer equal to that of the answer of the book.